Request help to show me the step by step instructions on how to solve Problem #1 below:
For each function,

Answer by COCOVALENZUELA(2)

find f(-2), f(-1/2), f(3) and f(7)
1. f(-2) = 5(-2) + 2
= -10 +2
= -8
2. f(-1/2)= 5(-1/2)+2
= - 2.5 + 2
= - 0.5
OK TAKE THESE AND WORK THE LAST ONE TO TEST YOUR COMPREHENSION